OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (−2, 0) and (2, 8)?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ 0}})\quad &({\color{red}{ 2}}\quad ,&{\color{blue}{ 8}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

oh k thank u

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

is the answer will be y = −2(x − 2)

OpenStudy (anonymous):

need help

OpenStudy (marissalovescats):

It depends on which point you chose to use. I'm guessing you chose (0,-2) ?

OpenStudy (marissalovescats):

Sorry (-2,0) lol

OpenStudy (marissalovescats):

Remember that two negatives equal a positive. Your slope would be 2 because: 8-0/2-(-2) = 8/4= 2 So if you used the point (-2,0) y-0=2(x-(-2) y-0=2(x+2) y=2x+4

OpenStudy (marissalovescats):

You could use the other point as well (2,8) y-8=2(x-2) y-8=2x-4 y=2x+4

OpenStudy (anonymous):

oh k this are the answers y = 2(x + 2) y = 2(x − 2) y = −2(x − 2) y = −2(x + 2)

OpenStudy (marissalovescats):

Well it would be y=2(x+2) because when you distribute the 2 you get y=2x+4

OpenStudy (anonymous):

oh k thank u

OpenStudy (marissalovescats):

You're welcome.